an energy source forces a constant current of 2.5amps for 20 seconds to flow through a light bulb. if 3kilojoules is given of in the form of light and heat energy calculate the charge of the current and the voltage drop across the bulb.


1
Expert's answer
2022-05-06T01:50:02-0400

The amount of charge

"q=It=2.5*20=50\:\rm C"

The voltage drop

"V=\frac{W}{q}=\frac{3000}{50}=60\:\rm V"